OpenStudy (anonymous):

A jogger desires to run a certain course in 1/4 less time than she usually takes; by what percent must she increase her average running speed to accomplish this goal?

OpenStudy (anonymous):

In movement problem, the speed is inversly proportional to the time of object moving. v1 is regular speed and responding time is t1 v2 is a goal-speed and responding time is t2 t1=4, t2 is less than t1 1 unit, t2=3 v1/v2 = t2/t1 = 3/4 v2=4/3 *v1 Percent increase 33 1/3%